OpenStudy (anonymous):

u substitution x^4-8x^2-9 = 0 u = x^2 u²-8u-9 = 0 can this be factored or do I need the quadratic?

jimthompson5910 (jim_thompson5910):

Two numbers that multiply to -9 and add to -8: -9 and 1 So replace -8u with -9u+1u to get u²-8u-9 = 0 u²-9u+1u-9 = 0 u(u-9)+1(u-9)=0 (u+1)(u-9)=0 u+1=0 or u-9=0 u=-1 or u=9 Since u = x^2, we know that x^2=-1 or x^2=9 So x = i, x=-i, x=3, or x=-3

OpenStudy (anonymous):

holy pellet I can't believe I didn't see that lmao. thanks jim. I must be tired.

OpenStudy (anonymous):

u²-8u-9 = 0 u²-9u+u-9 = 0 (u-9)(u+1)=0 u=9,u=-1 u=x^2 x^2=9, x^2=-1 x=3 x=-3 ,

jimthompson5910 (jim_thompson5910):

lol the solution is imaginary so that may explain why you didn't see it

OpenStudy (anonymous):

no i didn't see 9-1 = 8 lol.

OpenStudy (anonymous):

i was like, derp 3 x 3 does not = 8 when I add them.. fudge..